For the vectors π‘Ÿβƒ—1 = 2π‘–Μ‚βˆ’ 3 𝑗̂+ 4π‘˜Μ‚ π‘Žπ‘›π‘‘ π‘Ÿ2Μ‚ = 4𝑖̂+ 2π‘˜Μ‚


(a) Find the angle between them.


(b) Find the cross-product of the vectors.

1
Expert's answer
2021-11-15T17:54:44-0500

(a) Find the dot product and magnitudes:

"\vec r_1·\vec r_2=2·4+(-3)·2=2.\\
|\vec r_1|=\sqrt{2^2+(-3)^2}=3.6\\
|\vec r_2|=\sqrt{4^2+2^2}=4.5."

Find the angle between them:


"\cos\theta=\frac{\vec r_1·\vec r_2}{|\vec r_1|·|\vec r_2|}=\frac{2}{3.6·4.5}=0.12,\\\space\\
\theta=83.1°"

(b) Find the cross-product of the vectors:

"\vec r_1\times\vec r_2=|\vec r_1|·|\vec r_2|\sin\theta=1.95"